Annual demand for the Dobbs model airplane kit is 80,000 units. Albert Dobbs, president of Dobbs' Terrific Toys, controls one of the largest toy companies in Nevada. He estimates that the ordering cost is $40 per order. The carrying cost is $7 per unit per year. It is 25 days from the time that Albert places an order for the model airplane kits until they are received at his warehouse. During this time, the daily demand is estimated to be 450 units.

(a) Compute the EOQ, ROP, and optimal number of orders per year.

(b) Albert now believes that the carrying cost may be as high as $14 per unit per year. Furthermore, he estimates that the lead time may be 35 days instead of 25 days. Redo part (a), using these revised estimates.

Answer both parts especially part b
Answer #1

a)

Annual demand = 80000 = D
EOQ = Square root ((2*D*S)/h) S = ordering cost per order = $40 and h = holding cost = $7
EOQ = Sqrt((2*80000*40)/7) = 956.1828875

ROP = lead time* demand-EOQ = 25*450-956.1828875 = 10293.81711 [as lead time is greater than time between orders]

number of orders per year = Annual demand /EOQ = 80000/956.1828875 = 83.66600265

b)

Now, EOQ = Sqrt((2*80000*40)/14) = 676.1234038

ROP = lead time* demand-EOQ = 35*450-676.1234038= 15073.8766 [as lead time is greater than time between orders]

number of orders per year = Annual demand /EOQ = 80000/676.1234038 = 118.3215957